Show that energy transferred by energetic electron to helium atom by elastic collision is less than 0.1%.

1
Expert's answer
2021-12-20T10:29:07-0500

"\text{Take the frame of reference at a fixed helium nucleus}"

"m_1 - \text{helium atom }"

"m_2 - \text{electron}"

"v_1 = 0"

"EK = \frac{m_2v_2^2}{2}"

"m_2\vec{v_2}=m_1\vec{u_1} +m_2\vec{u_2}"

"m_1\vec{u_1} = m_2\vec{v_2}-m_2\vec{u_2}"

"max (\vec u_{1})= \frac{2\vec v_2m_2}{m_1}"


"\Delta EK_{m_1}= \frac{m_1u_1^2}{2}"

"max(\Delta EK_{m_1})= \frac{2m_2^2v_2^2}{m_1}"

"\text{maxpercentage}= \frac{max(\Delta EK_{m_1})}{EK}*100= \frac{4m_2}{m_1}*100"

"\text{The helium nucleus - is made up of two protons and two neutrons.}"

"\text{The mass of a proton is 1836 heavier than the mass of an electron}"

"\text{maxpercentage}= = \frac{4m_2}{4*1836*m_2}*100= 0.054 \%"

"\text{maxpercentage}<0.1 \%"


"\text{Proven}"